OpenStudy (anonymous):

What is the equation of a line that passes through (4, 7) and has a slope of −14? y=x4−74 y=−x4+74 y=−x4+8 y=x4+8 all i need is the answer

OpenStudy (anonymous):

um could you check if you wrote the question correctly really quick?

OpenStudy (anonymous):

it should be -1/4 as a fraction

OpenStudy (anonymous):

What is the equation of a line that passes through (4, 7) and has a slope of −1/4? y=x/4−7/4 y=−x/4+74 y=−x/4+8 y=x/4+8 all i need is the answer

OpenStudy (anonymous):

I'm getting the 3rd one

OpenStudy (anonymous):

i guessed the 2nd,the 3rd was right